OpenStudy (anonymous):

Solve the following system of equations algebraically: 2x+3y=11 5x-2y=-20 The solution is x=,y=

OpenStudy (anonymous):

Solve for x or y and substitute into the other equation.

OpenStudy (anonymous):

can you show me an example?

OpenStudy (anonymous):

For example 3x + 2y = 5 4x +5y=10 From first equation 3x=5-2y x = (5-2y)/3 Substitute x value in other equation 4[(5-2y)/3]+5y=10 You have only y values. Solve for y

OpenStudy (anonymous):

ehhhhh this is confusing.... this iswhy i dont think im takin the math regents..